OpenStudy (anonymous):

solve & simplify radicals: 115=10x^2-5

OpenStudy (saifoo.khan):

\[115 = 10x^2 - 5\] \[120 = 10x^2\] \[x^2 = \frac{120}{10}\] \[x^2 = 12\]

OpenStudy (saifoo.khan):

Now, \(x = \sqrt{12 }\) and \(x = -\sqrt{12}\)

OpenStudy (saifoo.khan):

\(x = - 2 \sqrt3 \), \(x = 2 \sqrt3\)